OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? a. y = 2x + 1 b. y = 2x + 3 c. y = 2x - 8 d. y = 2x - 3

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\] this is the point slope formula m=slope (x1,y1)

pooja195 (pooja195):

\[\huge~\rm~y-5=2(x-2)\] Can you try to solve for y?

OpenStudy (anonymous):

well the second equation would end up being y-5= 2x-4?